mead. If you were to make a random copolymer consisting of 30% Cand 70% G and subject it to in vitro translation, what amino acids would be incorporated into polypeptides? What percentage of the total would each represent? (10 pts) Below is a schematic representation of one VNTR studied in 20 individuals Since it is a single locus, each person has two nt the two homologou chromosomes There are five different (labeled A-E
Answer #1

So, the entire sequence contains C and G bases only.

Possible codons are CCC, CCG, CGC, CGG, GCC, GCG, GGC and GGG.

As per genetic code.

CCC and CCG code for Proline

CGC and CGG code for Arginine

GCC and GCG code for Alanine

GGC and GGG code for glycine.

So, these are the four amino acids incorporated into polypeptide chain.

Percentage calculation:

C and G are in 30% and 70% proportion.

So, the probabilities of a base being C and G are 0.3 and 0.7, respectively.

CCC and CCG code for Proline.

Probability of finding CCC codon is 0.3 x 0.3 x 0.3 = 0.027

Probability of finding CCG codon is 0.3 x 0.3 x 0.7 = 0.063

Probability of finding proline codon is 0.027 + 0.063 = 0.09

CGC and CGG code for Arginine.

Probability of finding CGC codon is 0.3 x 0.7 x 0.3 = 0.063

Probability of finding CGG codon is 0.3 x 0.7 x 0.7 = 0.147

Probability of finding arginine codon is 0.063 + 0.147 = 0.21

GCC and GCG code for Alanine

Probability of finding GCC codon is 0.7 x 0.3 x 0.3 = 0.063

Probability of finding GCG codon is 0.7 x 0.3 x 0.7 = 0.147

Probability of finding alanine codon is 0.063 + 0.147 = 0.21

GGC and GGG code for glycine

Probability of finding GGC codon is 0.7 x 0.7 x 0.3 = 0.147

Probability of finding GGG codon is 0.7 x 0.7 x 0.7 = 0.343

Probability of finding glycine codon is 0.147 + 0.147 = 0.49

Sum of probabilities of finding each amino acid should be 1.

0.09+0.21+0.21+0.49 = 1